what is the formula for a rectangle with a perimeter of 54"? its width is 1/2 the legnth.

Answer by checkley77(12844)

W=L/2
2L+2W=54
2L+2(L/2)=54
2L+L=54
3L=54
L=54/3
L=18 ANS. FOR THE LENGTH.
W=18/2=9 ANS. FOR THE LENGTH.
PROOF:
2*9+2*18=54
18+36=54
54=54